# Wall-resolved large eddy simulations of a pitching NACA0012 incurring in deep dynamic stall

### v1

@@ 2.2 Numerical methods
The steady incompressible RANS equations and the filtered unsteady NS ones are solved using ANSYS Fluent 2023R2 (ANSYS, 2023). A second-order upwind discretization in space is adopted. Gradients are retrieved through a least square cell-based method, and fluxes are obtained with the Rhie-Chow momentum-based formulation. The SIMPLE method is leveraged to solve the pressure-velocity equations in the steady framework while for the LES solution SIMPLEC method is used. The time evolution is obtained through a bounded second-order implicit time integration scheme.
Moving on to boundary conditions, the farfield marker present in the original mesh is split in two parts. For this reason, a marker is set for cells with positive $x$, corresponding to the coordinate around which the grid rotates and being $x$ the direction of the free-stream flow. Farfield faces are thus separated using the mark imposing a velocity inlet and a pressure outlet boundary conditions. The inflow velocity is imposed together with the turbulence intensity of 0.08% as in the wind tunnel. A non-slip and non penetrating condition is imposed on the airfoil. The surfaces at the tips of the extruded O-grid features a translational conformal periodic condition to grant a nominally infinite wing.
RANS system is closed using the $k-\omega$ SST turbulence model with the addition of the $\gamma$ intermittency equation (Menter et al., 2004), in which a differential transport equation is solved for the intermittency, to better describe the transition of the boundary layer from laminar to turbulent.
@@
The airfoil pitching motion is prescribed through the rigid motion of the entire grid. Indeed, the angular velocity is imposed in Fluent using a user-defined expression equal to $\dot{\alpha}(t)=\omega\ \alpha_{s}\cos(\omega t)$. The mean rotation, instead, is granted through a rigid rotation of the mesh before starting the simulation. In order to speed up temporal convergence, the simulation is performed in three different stages. Firstly, a steady RANS simulation with the mesh rotated at $\alpha_{0}$ is performed until force coefficient convergence is reached. After that, the time-dependent pitching simulation is evolved. The mesh rotation is imposed to obtain a negative pitch rate after the steady simulation (Baldan and Guardone, 2023). This allows to reach unsteady convergence faster since the flow is evolved at lower angle of attack and is attached to the profile. The first pitch down quarter of the cycle is neglected. Finally, an entire pitching cycle is computed, and data are collected.
Following each simulation, the pressure coefficient $C_{p}$ and skin friction coefficient $C_{f}$ distributions over the wing section are extracted every 40 time steps. These unstructured outputs from Fluent are subsequently mapped onto the structured mesh of the wing section using a k-d tree algorithm, which efficiently identifies the nearest neighbors Virtanen et al. (2020). For post-processing, the skin friction value at each point is converted to its signed magnitude. The sign is determined based on the direction calculated from the $x$ and $y$ components of $C_{f}$. The positive geometric direction is aligned with the curvilinear coordinate that extends from the leading edge to the trailing edge. When the geometric direction and the numerically computed $C_{f}$ direction coincide, the sign is positive. Conversely, if they oppose each other, the sign is negative. This process allows for the identification of regions with flow separation.
@@ 3 Results
The present section discusses the results obtained with the previously described numerical models. Specifically, the effect of the span-to-chord ratio is investigated in Section 3.1, while Section 3.2 reports the comparison of the proposed LES with other numerical approaches.
3.1 Span-to-chord ratio effect
Figure 3: Spanwise size comparison of lift and drag coefficients. Experimental data from Lee and Gerontakos (2004).
$z/c=0.2$ Refer to caption
$z/c=0.4$ Refer to caption
$z/c=0.8$ Refer to caption
$z/c=1.2$ Refer to caption
Figure 4: Effect of spanwise extrusion on instantaneous velocity magnitude contour.
Figure 5: Effect of spanwise extrusion on instantaneous friction coefficient contour over the suction side of the airfoil.
@@ 3.1 Span-to-chord ratio effect
Figure 8: Q-criterion iso-surface with velocity magnitude contours at different instants of the pitching cycle. $\uparrow$ indicates the upstroke phase while $\downarrow$ the downstroke one. Mesh extends for $1.2c$ in spanwise direction.
The initial stage of the investigation centers on the impact of employing computational grids with limited spanwise dimensions when encountering dynamic stall phenomena. Four distinct grid sizes are employed for this analysis: $0.2c$, $0.4c$, $0.8c$, and $1.2c$.
@@
The primary focus is on the aerodynamic loads experienced throughout the complete pitching cycle. Figure 3 depicts the lift and drag coefficients as functions of the angle of incidence. When the flow remains predominantly attached, during both the initial portion of the upstroke and the last part of the downstroke, the behavior displayed by the grids exhibits a high degree of similarity. However, a significant divergence is observed between the two smaller grids when the LSB detaches and subsequently induces the DSV at an angle of incidence of approximately 22 degrees. Notably, the first peak in lift during the pitch down phase surpasses the values obtained with the other two grids. This disparity becomes even more pronounced at lower angles of attack, where the interaction between the TEV and the DSV results in a lift peak that even exceeds the value observed during the climb phase. A comparable trend is also observed with regards to the drag coefficient. Lastly, larger grids exhibit a trend similar to experimental data during the pitch down, although a lift offset is present.
@@
Figure 4 illustrates the instantaneous velocity magnitude contours at peculiar angles, providing a more detailed understanding of the influence of the span size. The figure initially confirms that mesh size has no impact on flow structures during the upstroke phase. This is evident from the qualitatively similar appearance of the subfigures in the first row for each grid. Focusing on the angle of maximum incidence where the DSV is fully developed, we observe a significant effect of the span-to-chord ratio on the flow field. For grids with a ratio of 0.2 and even 0.4, the boundary conditions enforce a predominantly two-dimensional character on the flow. This is manifested by a more defined DSV core and a higher peak velocity magnitude. In contrast, grids with larger spanwise extensions exhibit smaller turbulence scale structures, likely due to interactions in the spanwise direction. This phenomenon becomes particularly pronounced during the descent phase of the profile. Here, the smaller grids reveal the presence of a distinct vortical structure absent in the other cases. A peculiar difference is available at $\alpha$ equal to 10 degrees, where the two grids with larger spans exhibit distinct behaviors. The largest grid displays a completely stalled flow, while the grid with a span of $0.8c$ still presents oscillatory patterns in the wake region.
@@
Further insight into the spanwise effects is provided by analyzing the skin friction coefficient distribution on the wing section’s upper surface. Figure 5 presents the $C_{f}$ distribution for the previously investigated angles of attack across all four meshes. The observations regarding the upstroke phase are corroborated by the alignment of the recirculation zone, identified by negative $C_{f}$ values. The most significant discrepancies arise at the maximum angle of incidence, along with the descending phase at 20 and 10 degrees. These contours demonstrate the development of three-dimensionality within the flow field when sufficient spanwise extension is available. Interestingly, a departure from prior observations is seen at 22 degrees, where the $C_{f}$ trends exhibit a high degree of similarity between all meshes. Additionally, the qualitative behavior at 10 degrees appears comparable for the two largest grids.
@@
Figure 6 depicts the time-dependent evolution of the DSV and TEV by presenting the spanwise-averaged $C_{p}$ and $C_{f}$ trends on the suction side of the airfoil. This investigation offers additional insights beyond the previous analyses. We observe the boundary layer undergoing the transition from laminar to turbulent regime. The transition starts at the trailing edge around an angle of attack of 2 degrees and progresses towards the LSB at 10 degrees. Beyond 15 degrees, the development of the DSV becomes evident. It grows until it reaches the trailing edge at 22 degrees, where it interacts with the TEV. The subsequent behavior varies depending on the grid size. For grids with $z/c$ of 0.2 and 0.4, two distinct pressure coefficient peaks are observed at the trailing edge, followed by a weaker third one. Additionally, the detachment of another vortex from the mid-chord region is visible. In contrast, the behavior for the grids with larger spans is different. The trailing edge experiences a pressure coefficient peak, albeit lower in magnitude compared to the smaller grids, upon the first interaction with the DSV. This is followed by a weaker peak, and subsequently, the flow becomes completely stalled, exhibiting no discernible trends.
@@
To conclude the mesh analysis, Figure 7 presents the wall units averaged in the spanwise direction for the grid with the largest span extent, evaluated at the maximum angle of incidence, 25 degrees. The reported values adhere to the recommended practices for WRLES. However, the $\eta^{+}$ value is more conservative with respect to the just mentioned LES range and is closer to 1. This value satisfies the requirement imposed by the RANS turbulence model adopted in the flow initialization.
@@
Figure 8 presents a visualization of the three-dimensional dynamic stall structures using the Q-criterion iso-surface with velocity magnitude contours. A key observation is the difference in the fields at identical angles of attack but during ascent and descent phases. Analyzing each phase, the laminar-to-turbulent boundary layer transition is visible during the 10-degree in the upstroke phase, characterized by the presence of instabilities and hairpin structures. At an angle of approximately 20 degrees, the DSV begins to develop, exhibiting a primarily two-dimensional structure. As the angle of attack reaches its maximum, where the DSV is fully developed and interacts with the TEV, the flow is completely three-dimensional exhibiting spanwise variations. The descent phase at 20 and 10 degrees is still fully stalled and three-dimensional. Subsequently, as the angle of attack approaches zero, the flow starts to laminarize. Finally, at the lowest incidence, the flow presents some instabilities on the lower side of the wing section.
@@ 3.2 Comparison with RANS and hybrid RANS/LES approaches
In conclusion, we conduct a comparative analysis of the just presented LES results with those obtained from previous studies by Baldan and Guardone (2024b). Two approaches were employed in the prior work: a two-dimensional unsteady RANS simulation and a three-dimensional approach utilizing hybrid RANS/LES techniques. For the RANS simulation, the $k-\omega$ SST turbulence model was implemented in conjunction with the $\gamma$ intermittency equation, referred to as iSST, to enhance the description of the laminar-to-turbulent transition within the boundary layer. The hybrid model employed a Stress-Blended Eddy Simulation approach coupled with the same RANS turbulence model. The reference data for the LES simulations is based on the mesh with a span-to-chord ratio of 1.2.
Figure 9 presents a comparison of the profile loads obtained from the numerical simulations and the experimental data by Lee and Gerontakos (2004). The results reveal a high degree of agreement between the numerical models for attached flow conditions, with their curves nearly overlapping. Notably, the most significant discrepancy between the numerical and experimental data lies in the angle of attack corresponding to peak lift. Interestingly, the simulations anticipate, with respect to the experimental data, the angle at which the DSV is convected downwards and this behavior is consistent across all numerical models, including the WRLES approach. In contrast, during the initial descent phase of the profile, characterized by fully stalled flow, the models exhibit some divergence. However, as the angle of incidence decreases, they converge towards a unified solution. It has to be highlighted that in the stall portion cycle-to-cycle variations are not considered due to the large computational resource requirements of large eddy simulations. Despite an offset from the experimental data persists, the overall behavior of simulations aligns well with the experiments.
@@ 4 Conclusions
This work employed wall-resolved large eddy simulations to investigate the impact of grid resolution on pitching airfoils. The study revealed a critical dependency of key flow structures, particularly the three-dimensionality of the dynamic stall vortex, on the spanwise grid size. While existing literature offers case studies with limited span-to-chord ratios, 0.1-0.2, these were found to be inadequate for capturing the full complexity of the phenomenon. Consequently, results obtained with such grids are unreliable except for the attached flow region involving the laminar separation bubble formation and the dynamic stall vortex generation. This highlights the growing challenge of studying the phenomenon with high-fidelity models at increasing Reynolds numbers due to the ever-increasing mesh size requirements.
Furthermore, a comparison with computational models based on RANS equations suggested a surprising result. The key discrepancies between numerical simulations and experimental data are not primarily due to turbulence models, as commonly reported. These discrepancies might instead be linked to other factors such as wind tunnel effects or the inherent limitations in accurately replicating the surface roughness of the wind tunnel wing section model within the computational models.

### v2

@@ 1.2 Numerical methods
The steady incompressible RANS equations and the LES ones are solved using ANSYS Fluent 2023R2 (ANSYS, 2023). A second-order upwind discretization in space is adopted. Gradients are retrieved through a least square cell-based method, and fluxes are obtained with the Rhie-Chow momentum-based formulation. The SIMPLE method is used to solve the pressure-velocity equations in the steady framework. The SIMPLEC method is used to compute the LES solution. The time evolution is obtained through a bounded second-order implicit time integration scheme.
Moving on to boundary conditions, the farfield tag present in the original mesh is split into two parts during the simulation evolution. This is necessary since the mesh rigidly rotates and it is not possible to keep the same boundary condition definition. For this reason, a marker is set for cells with positive $x$, corresponding to the coordinate around which the grid rotates and being $x$, the direction of the free-stream flow. A velocity inlet condition is imposed to farfield faces associated to the mark, $x<0$, while a pressure outlet boundary condition is set for the others, $x\geq 0$.
At the farfield, inflow velocity is imposed together with the turbulence intensity of 0.08% as in the wind tunnel. A non-slip and non penetrating condition is imposed on the airfoil. The surfaces at the tips of the extruded O-grid features a translational conformal periodic condition to grant a nominally infinite wing.
RANS system is closed using the $k-\omega$ SST turbulence model with the addition of the $\gamma$ intermittency model (Menter et al., 2004), in which a differential transport equation is solved for the intermittency, to better describe the transition of the boundary layer from laminar to turbulent.
@@
The airfoil pitching motion is prescribed through the rigid motion of the entire grid. The angular velocity is imposed using a user-defined expression equal to $\dot{\alpha}(t)=\omega\ \alpha_{s}\cos(\omega t)$. The mean angle of attack $\alpha_{0}$ is imposed through a rigid rotation of the mesh before starting the simulation. In order to speed up temporal convergence, the simulation is performed in three different stages. Firstly, a steady RANS simulation with the mesh rotated at $\alpha_{0}$ is performed until force coefficient convergence is reached. Then, the time-dependent pitching simulation is evolved. The mesh rotation is imposed to obtain a negative pitch rate after the steady simulation (Baldan and Guardone, 2023). The first pitch down quarter of the cycle is neglected. Finally, an entire pitching cycle is computed, and data are recorded.
Following each simulation, the pressure coefficient $C_{p}$ and skin friction coefficient $C_{f}$ distributions over the wing section are extracted every 40 time steps and are mapped onto the structured mesh of the wing section using a k-d tree algorithm, which efficiently identifies the nearest neighbors. For post-processing, the skin friction value at each point is converted to its signed magnitude. The sign is determined based on the direction calculated from the $x$ and $y$ components of $C_{f}$. The positive geometric direction is aligned with the curvilinear coordinate that extends from the leading edge to the trailing edge. If the geometric direction and the numerically computed $C_{f}$ direction coincide, the sign is positive. Conversely, if they oppose each other, the sign is negative. This process allows for the identification of regions with flow separation.
@@ 2 Results
The present section discusses the results obtained with the numerical models described previously. Specifically, the effect of the span-to-chord ratio is investigated in Section 2.1, while Section 2.2 reports the comparison of the proposed LES with other numerical approaches.
2.1 Span-to-chord ratio dependence
Figure 3: Spanwise size comparison of lift and drag coefficients for the simulated cycle. Experimental data from Lee and Gerontakos (2004).
Figure 4: Effect of span-to-chord ratio on instantaneous velocity magnitude contour at mid-section.
Figure 5: Effect of span-to-chord ratio on instantaneous friction coefficient contour over the suction side of the airfoil.
@@ 2.1 Span-to-chord ratio dependence
Figure 8: Q-criterion iso-surface with velocity magnitude contours at different instants of the pitching cycle. $\uparrow$ indicates the upstroke phase while $\downarrow$ the downstroke one. The model extends for $1.2c$ in spanwise direction.
The initial stage of the investigation centers on the impact of employing wing geometries with limited span-wise dimensions to simulate dynamic stall phenomena. Four different span-to-chord ratios are used: $L_{z}/c=0.2,0.4,0.8,1.2$.
@@
The primary focus is on the aerodynamic loads experienced throughout the complete pitching cycle. Figure 3 depicts the instantaneous lift and drag coefficients as functions of the angle of incidence. If the flow remains predominantly attached, the behavior displayed at different span-to-chord ratios exhibits a high degree of similarity. During the initial portion of the upstroke, ranging from $\alpha=-5^{\circ}$ to $\alpha\approx 22^{\circ}$, the maximum difference of $C_{L}$ value is equal to 0.05. However, a significant divergence is observed between the two smaller span-to-chord ratios when the LSB detaches and subsequently induces the DSV at an angle of incidence of approximately 22 degrees. Notably, the first peak in lift during the pitch-down phase exceeds the values obtained with the larger span-to-chord ratios. This disparity becomes even more pronounced at lower angles of attack, where the interaction between the TEV and the DSV results in a lift peak that even exceeds of 0.5 the $C_{L}$ value observed during the climb phase and of 0.3 the $C_{D}$ one. Lastly, larger grids exhibit a trend similar to the experimental data during pitch-down, although a $\approx 0.25\ C_{L}$ offset is present.
@@
Figure 4 illustrates the magnitude contours of instantaneous velocity at peculiar angles, providing a more detailed understanding of the influence of the span size. The figure initially confirms that span-to-chord ratio has no impact on flow structures during the upstroke phase. This is evident from the qualitatively similar appearance of the subfigures in the first row for each span-to-chord ratios. Focusing on the angle of maximum incidence where the DSV is fully developed, we observe a significant effect of the span-to-chord ratio on the flow field. For geometries with a ratio of 0.2 and even 0.4, the boundary conditions enforce a predominantly two-dimensional character on the flow. This is manifested by a more defined DSV core and a higher peak velocity magnitude. In contrast, geometries with larger spanwise extensions exhibit smaller turbulence scale structures, likely due to interactions in the spanwise direction. This phenomenon becomes particularly pronounced during the descent phase of the airfoil. Here, the geometries with smaller span-to-chord ratio exhibit a distinct vortical structure absent in the other cases. A peculiar difference is seen at $\alpha$ equal to 10 degrees, where the two geometries with larger spans exhibit different behaviors. The geometry with the largest span-to-chord ratio displays a completely stalled flow, whereas the geometry with a span of $0.8c$ still presents oscillatory patterns in the wake region.
@@
Further insight into the span-to-chord ratio effect is provided by analyzing the skin friction coefficient distribution on the wing section’s upper surface. Figure 5 presents the $C_{f}$ distribution for the previously investigated angles of attack across all four meshes. The observations regarding the upstroke phase are corroborated by the alignment of the recirculation zone, identified by negative $C_{f}$ values. The most significant discrepancies arise at the maximum angle of incidence, along with the descending phase at 20 and 10 degrees. These contours demonstrate the development of three-dimensional structures within the flow field when sufficient spanwise extension is available. Interestingly, a departure from prior observations is seen at 22 degrees, where the $C_{f}$ trends exhibit a high degree of similarity between all geometries. Additionally, the qualitative behavior at 10 degrees appears comparable for the two geometries with the larger span-to-chord ratios.
@@
Figure 6 depicts the time-dependent evolution of the DSV and TEV by presenting the spanwise-averaged $C_{p}$ and $C_{f}$ trends on the suction side of the airfoil. The boundary layer undergoing the transition from laminar to turbulent regime. The transition starts at the trailing edge around an angle of attack of 2 degrees and progresses towards the LSB at 10 degrees. Beyond 15 degrees, the development of the DSV becomes evident. The DSV grows until it reaches the trailing edge at 22 degrees, where it interacts with the TEV. The subsequent behavior varies depending on the grid size. For geometries with $L_{z}/c$ of 0.2 and 0.4, two different pressure coefficient peaks are observed at the trailing edge, followed by a weaker third one. Additionally, the detachment of another vortex from the mid-chord region is visible. In contrast, the behavior for the geometries with larger spans is different. The trailing edge experiences a pressure coefficient peak, albeit lower in magnitude compared to the smaller span-to-chord ratios, upon the first interaction with the DSV. This is followed by a weaker peak, and subsequently, the flow becomes completely stalled, exhibiting no discernible trends.
@@
To assess the resolution of the computational mesh, Figure 7 presents the wall units averaged in the spanwise direction for the computational mesh of the geometry with the largest span extent, evaluated at the maximum angle of incidence, 25 degrees. The reported values adhere to the recommended practices for WRLES. However, the $\eta^{+}$ value is lower than that imposed by LES best practice and is closer to 1. This value satisfies the requirement imposed by the RANS turbulence model adopted in the flow initialization.
@@
Figure 8 presents a visualization of the three-dimensional dynamic stall structures using the Q-criterion iso-surface with velocity magnitude contours. A key observation is the difference in the fields at identical angles of attack during the ascent and descent phases. Analyzing each phase, the laminar-to-turbulent boundary layer transition is visible during the 10-degree in the upstroke phase, characterized by the presence of instabilities and hairpin structures. At an angle of approximately 20 degrees, the DSV begins to develop, exhibiting a primarily two-dimensional structure. As the angle of attack reaches its maximum, where the DSV is fully developed and interacts with the TEV, the flow is completely three-dimensional and exhibits spanwise structures. The descent phase at 20 and 10 degrees is still fully stalled and three-dimensional. Note that this portion of the pitching cycle is not simulated in previous studies of dynamic stall that use a ramp-up motion, but is critical for describing dynamic stall in its entirely. Furthermore, it is the condition when the largest differences occur depending on the model span size. At the end of the descent phase, as the angle of attack approaches zero, the flow starts to laminarize. Finally, at the lowest incidence, the flow presents some instabilities on the lower side of the wing section.
@@ 2.2 Comparison with RANS and hybrid RANS/LES approaches
We present now a comparative study of the LES results from the previous section with those obtained from previous studies by Baldan and Guardone (2024b). Two approaches were employed by  Baldan and Guardone (2024b): a two-dimensional unsteady RANS simulation and a three-dimensional approach utilizing hybrid RANS/LES techniques. For the RANS simulation, the $k-\omega$ SST turbulence model was implemented in conjunction with the $\gamma$ intermittency model, referred to as iSST, to enhance the description of the laminar-to-turbulent transition within the boundary layer. The hybrid model employed a Stress-Blended Eddy Simulation approach coupled with the same RANS turbulence model. The reference data for the LES simulations is based on the mesh with a span-to-chord ratio of 1.2.
Figure 9 presents a comparison of the profile loads obtained from the numerical simulations and the experimental data by Lee and Gerontakos (2004). The $C_{L}$ results present a maximum 0.1 difference between the numerical models for attached flow conditions, with their curves nearly overlapping. Notably, the most significant discrepancy between the numerical and experimental data lies in the angle of attack corresponding to peak lift which is $22^{\circ}$ and $24^{\circ}$, respectively. Interestingly, the simulations anticipate, with respect to the experimental data, the angle at which the DSV is convected downwards and this behavior is consistent across all numerical models, including the WRLES approach. In contrast, during the initial descent phase of the profile $\alpha=24^{\circ}-18^{\circ}$, characterized by fully stalled flow, the models exhibit some divergence. However, as the angle of incidence decreases, they converge towards the same solution. Note that in the stall portion cycle-to-cycle variations are not considered due to the large computational resource requirements of large eddy simulations. Despite a $C_{L}$ offset equal to $0.2$ from the experimental data persists, the overall behavior of simulations aligns well with the experiments.
@@ 3 Conclusions
This work investigated the impact of the span-to-chord ratio on pitching airfoils using wall-resolved large eddy simulations. The study revealed a critical dependency of key flow structures, particularly the three-dimensionality of the dynamic stall vortex, on the spanwise size. While existing literature reports case studies with limited span-to-chord ratios, from 0.1 up to 0.2, these were found not adequate to capture the full complexity of the phenomenon. Consequently, results obtained with such geometries are can be accepted only in the attached flow region, typical of ramp-up motions, involving the laminar separation bubble formation and the dynamic stall vortex generation. On the contrary, a complete pitching motion was used in the present work to highlight these limitations. These results point to the growing challenge of studying the phenomenon with high-fidelity models at increasing Reynolds numbers due to the ever-increasing mesh size requirements.
Furthermore, a comparison with computational models based on RANS equations suggested a surprising result. The key discrepancies between numerical simulations and experimental data are not primarily due to turbulence models, as commonly reported. These discrepancies seem instead to be linked to other factors, such as possibly wind tunnel effects or the inherent limitations in accurately replicating the surface roughness of the wind tunnel wing section model within the computational models.